# PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems

### v4

@@ Abstract.
Generative retrieval methods utilize generative sequential modeling techniques, such as transformers, to generate candidate items for recommender systems. These methods have demonstrated promising results in academic benchmarks, surpassing traditional retrieval models like two-tower architectures. However, current generative retrieval methods lack the scalability required for industrial recommender systems, and they are insufficiently flexible to satisfy the multiple metric requirements of modern systems.
This paper introduces PinRec, a novel generative retrieval model developed for applications at Pinterest. PinRec utilizes outcome-conditioned generation, enabling modelers to specify how to balance various outcome metrics, such as the number of saves and clicks, to effectively align with business goals and user exploration. Additionally, PinRec incorporates multi-token generation to enhance output diversity while optimizing generation. Our experiments demonstrate that PinRec can successfully balance performance, diversity, and efficiency, delivering a significant positive impact such as +2% sitewide clicks and +4% search repins. This paper marks a significant milestone in generative retrieval, as it presents, to our knowledge, the first rigorous study on implementing generative retrieval at the scale of Pinterest.
generative retrieval, recommender systems, transformer, item recommendation
@@
Figure 1. Illustration of PinRec, a generative item retrieval technique for heterogeneous user journeys on Pinterest. Sequences of user searches, engagements, and outcome-conditioning (bottom) are used to recommend Pins (top).
@@ 1. Introduction
Pinterest is a visual discovery platform where over 550 million monthly active users search for, save, and shop for ideas, represented by Pins (pins). Given the massive volume of Pins — for instance, users save over 1.5 billion Pins per week (pins2) — it is essential to provide personalized recommendations in Pinterest feeds and search results. To achieve this, recommender systems often leverage a two-stage approach: retrieval and ranking. Retrieval models select thousands of potentially relevant candidate items, then ranking models prioritize the top items to display to the user. Since ranking models depend on the retrieval results, it is important for retrieval models to identify diverse, fresh, and relevant items.
@@
Traditionally, retrieval methods employ “two-tower” models. The query tower computes query embeddings, while the item tower computes item embeddings. These models retrieve items with the highest similarity scores given a query embedding. As an alternative to two-tower models, recent work has proposed generative retrieval. Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (vaswani2017attention), and then use those tokens to retrieve items. Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (KangM18; Recformer23; UnifyingGenRetrieval24).
Using generative retrieval models for large scale recommender systems poses many challenges at the scale of Pinterest. Firstly, the generative model should be capable of efficiently optimizing for different engagement outcomes such as maximizing social actions like saves and comments versus passive engagements like clicks, because business needs are constantly evolving to optimize combinations of different metrics. Secondly, the generated item sequences must possess sufficient diversity to facilitate users’ exploration of varied ideas; one-step generation of a single item representation may not suffice. Thirdly, given their increased complexity and sequential generation, generating a set of recommended items requires higher computational cost than computing an embedding with two-tower models.
To our knowledge, there has been no large scale generative retrieval model that has addressed all of the aforementioned challenges. Most methods developed within academic settings (KangM18; UniS4Rec22; Recformer23) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic. While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (HSTU24; Tiger23) and are not capable of optimizing multiple target outcomes. Moreover, these methods do not address the challenge of generating a sequence of diverse items. Because generating multiple items sequentially is computationally expensive, existing approaches typically generate only the next item rather than a full set at once.
Figure 2. Modeling diagram of PinRec
In this paper, we present PinRec, a generative retrieval technique for real-world recommender systems (Figure 2). PinRec builds on other generative retrieval models, leveraging transformers to perform next token prediction (KangM18; HSTU24; Tiger23). Unlike previous methods, however, PinRec addresses and resolves three aforementioned challenges for generation of optimal recommendations: multi-objective generative capabilities, computational efficiency, and item diversity. Our main contributions in the paper are summarized below.
@@
  • Outcome-Conditioned Generative Retrieval: PinRec employs outcome-conditioned generation, enabling the optimization of multiple business metrics. For each outcome (e.g., business metrics such as user saves or clicks), PinRec generates the next item token that is most likely to maximize that specific outcome, and then retrieves candidate items to balance these objectives according to the desired trade-offs (Fig. 2). Through both offline and online evaluations, we show that tailoring item recommendations to particular outcomes enables designers to effectively achieve those specified objectives.
  • Multi-Token Generation for Efficient Sequence Generation: One obstacle to generating an entire sequence of items for retrieval is that traditional models typically produce only one item at a time. To address this, we propose a multi-token generation framework (Multitoken24), in which PinRec generates multiple future items in parallel, as illustrated in Figure 2. This approach significantly reduces the number of steps needed to produce diverse output embeddings, resulting in substantial offline and online improvements in performance, diversity, and efficiency.
  • Efficient Serving: We detail how we optimize autoregressive generation for optimal item recommendation, implement the serving infrastructure, and allow for horizontal scaling to serve hundreds of millions of users. To our knowledge, this study is the first to disclose the details of infrastructure optimization for industry-scale generative retrieval systems.
@@ 2. Related Work
Sequential Recommendation Systems: Sequential recommendation systems operate on a sequences of items with which a user has interacted (KangM18). Earlier work in academia has used recurrent neural networks (RNN) (GRU4REC16), but later work has often adopted transformers (KangM18; BERT4REC19). Recent studies have harnessed the extensive knowledge of pre-trained language models (UniS4Rec22; Recformer23).
In industrial applications, a common approach is to apply sequential encoders to learn user representations (Pinnerformer22). The user representation can then be used in traditional recommender systems (e.g., embedding-based retrieval or ranking models) (Spotify20; rethink; Transact23; TWINV224). Another approach in sequential recommendation systems is the generative approach, where the model learns to generate the items to recommend. Tiger23 used transformers to generate items by producing a sequence of semantic ID tokens (RQVAE22; SemanticIDGeneral24). HSTU24 also proposed a generative retrieval method. However, these studies have mainly focused on generating a single item to optimize a single action. In contrast, our research aims to generate sequences of multiple items to optimize for multiple outcomes.
Multi-Objective Generation: We discuss generative models that consider multiple objectives. While RL algorithms have shown this capability in industry (SelfsupervisedRL20), studies in language modeling suggest RL often needs extensive tuning for optimal generation (PPO22). Recent research has introduced more streamlined and scalable techniques based on supervised learning, such as goal-conditioned methods (IterativeSupervised21) and prompting strategies (PromptingRL22). PinRec builds on these approaches to enhance scalability.
@@ 3. Problem Setup
User History. For each user $u$, we maintain their interaction history $H(u,t_{\text{max}})=\langle(i_{1},a_{1},s_{1},t_{1}),(i_{2},a_{2},s_{2},t_{2}),\ldots,(i_{m},a_{m},s_{m},t_{m})\rangle$, where each tuple $(i_{j},a_{j},s_{j},t_{j})$ represents an interaction with item $i_{j}\in\mathcal{I}$ via action $a_{j}\in\mathcal{A}$ on surface $s_{j}\in\mathcal{S}$ at timestep $t_{j}\leq t_{\text{max}}$. The sequence is chronologically ordered such that $t_{1}\leq t_{2}\leq\ldots\leq t_{m}$.
@@ 4. PinRec
In this section, we describe the model architecture and serving infrastructure for PinRec. We start with an overview of the model, including outcome-conditioned generation and multi-token prediction. Next, we explain how items in the sequence are encoded and processed, and how the training loss is computed. Finally, we describe inference mode operation and provide details about our real-time serving architecture, which is powered by Nvidia Triton.
@@ 4.1. Model Architecture
Outcome-Conditioned Generation. Unlike language modeling, where the output sequence receives a single reward value (PPO22), users interact with recommended items through diverse engagement types—saving, sharing, commenting, or clicking. To effectively optimize across these varied engagement metrics, generative retrieval methods must account for this behavioral diversity (MultiBehaviorSASRec24). Furthermore, practical deployment often requires specifying the target surface or context for the next recommendation, even when the user’s historical sequence spans multiple surfaces.
@@
Temporal Multi-token Prediction. Multi-token prediction in language models involves predicting multiple successive tokens at positions $t+1$ to $t+k$ given the current position $t$ (gloeckle2024better). This is typically implemented using separate output heads for each future position. However, in recommender systems, previous research (Pinnerformer22) suggests that the difference in time between actions is much more important than the position index alone, because a single position increment can correspond to very different time intervals for active users versus dormant users. To address this, we propose a time-based multi-token approach, in which tokens are generated for each time interval $\delta$ (e.g., generating recommendations for the user for the present, 30 seconds later, 1 minute later, and so on). Specifically, we condition our output head on the relative temporal distance from the current action, enabling a single parameterized head to generate predictions at various future time intervals:
(2) $\hat{\mathbf{i}}_{u,t+\delta}=O(\mathbf{h}_{u,t},\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k},\mathbf{e}_{\delta})$
where $\mathbf{e}_{\delta}$ is the learnable embedding for temporal offset $\delta\in\{1,2,\ldots,k\}$, and $\hat{\mathbf{i}}_{u,t+\delta}$ represents the predicted item at timestep $t+\delta$.
@@ 4.2. Input Sequence Representation
Given the heterogeneous nature of user interactions $(i_{j},a_{j},s_{j},t_{j})$ in our interaction history $H(u,t_{\text{max}})$, we need to transform each element into appropriate representations for the transformer architecture. A crucial design decision is defining representations for heterogeneous item types. There are two main ways to represent items in a sequence: (1) by using dense embedding vectors (HSTU24; KangM18), and (2) by learning a sequence of discrete IDs for each item, known as Semantic IDs (Tiger23; SemanticIDGeneral24). We explored both approaches and ultimately chose to use dense embedding vectors. In our experiments, we observed that Semantic IDs frequently suffered from representational collapse—many items were assigned the same Semantic ID, as also noted in (MultiBehaviorSASRec24; Hourglass24)—and consistently produced lower evaluation metrics (see Table 1(UnifyingGenRetrieval24). While optimizing performance with Semantic IDs remains an interesting direction for future work, our core contributions are representation-agnostic.
@@
Pin Embeddings. For Pins, we combine semantic embedding features with learned Id embeddings (10.1145/3640457.3688053). The semantic representation uses the pre-trained OmniSage embedding (badrinath2025omnisage), which captures visual and textual content along with the engagement via the Pin-board graph. Additionally, we implement a large-scale ID embedding system where each Pin identifier is hashed into $k$ different hash functions, each mapping to embedding tables of size $d/k$. These $k$ embeddings are concatenated to form a $d$-dimensional ID embedding, providing a distributed representation that scales efficiently. The ID embeddings constitute sparse parameters that allow us to scale model capacity with minimal impact on serving latency and computational cost, as only a small subset of embeddings are accessed for any given batch of items. The concatenated OmniSage and ID embeddings are passed through 2-3 fully connected layers with ReLU activation and layer normalization, followed by $L_{2}$ normalization to produce the final Pin embedding $\mathbf{i}_{\text{pin}}$.
Query Embeddings. For search queries, we use only semantic representations as ID embeddings have not provided value in offline metrics for this entity type. We leverage the pre-trained OmniSearchSage query representation (agarwal2024omnisearchsage) which captures semantic query understanding. This representation is passed through a similar MLP architecture (2-3 fully connected layers with ReLU activation and layer normalization) followed by $L_{2}$ normalization to generate the final query embedding $\mathbf{i}_{\text{query}}$.

### v5

# PinRec: Unified Generative Retrieval for Pinterest Recommender Systems

@@ Abstract.
Generative retrieval methods employ sequential modeling techniques, like transformers, to generate candidate items for recommender systems. These methods have demonstrated promising results in academic benchmarks, surpassing traditional retrieval models such as two-tower architectures. However, a key limitation is that current approaches require a separate model for each product surface, as building a unified model that accommodates the different business needs of various surfaces has proven challenging. Furthermore, existing methods often fail to capture the evolution of user interests over a sequence, focusing instead on only predicting the next item.
This paper introduces PinRec, a novel unified generative retrieval model for all of Pinterest’s recommendation surfaces, including home feed, search, and related pins. PinRec is pretrained on user activity sequences aggregated across surfaces, then fine-tuned for each surface using that surface’s impression data. This pretraining–fine-tuning approach enables a single unified model while still adapting to the needs of individual surfaces. To better align recommendations with surface-specific business goals, PinRec incorporates a novel outcome-conditioned generation mechanism that targets different outcomes for each surface, which further enhances the impact of fine-tuning. Our experiments show that PinRec balances performance, diversity, and efficiency, delivering significant gains such as +4% increase in search saves. To our knowledge, this paper presents the first rigorous study of a unified generative retrieval model built and deployed at Pinterest scale, marking a significant milestone in the field.
generative retrieval, recommender systems
@@
Figure 1. Illustration of PinRec, a unified generative retrieval model for heterogeneous user journeys on Pinterest. Sequences of user searches and activities are used to recommend Pins (top). PinRec powers retrieval in multiple surfaces optimizing surfaces’ unique business metrics like Saves for Home Feed and Product Click for Search
@@ 1. Introduction
Pinterest is a visual discovery platform where over 600 million monthly active users search for, save, and shop for ideas, represented by Pins (Pinterest, 2025). Given the massive volume of Pins—for instance, users save over 1.5 billion Pins per week (Pinterest, 2024)—it is essential to provide personalized recommendations in Pinterest feeds and search results. To achieve this, recommender systems often leverage a two-stage approach: retrieval and ranking. Retrieval models select thousands of potentially relevant candidate Pins, then ranking models prioritize the top Pins to display to the user. Since ranking models depend on the retrieval results, it is important for retrieval models to identify diverse, fresh, and relevant Pins.
@@
Traditionally, retrieval methods employ “two-tower” models. The query tower computes query embeddings, while the item tower computes Pin embeddings. These models retrieve Pins with the highest similarity scores given a query embedding. As an alternative to two-tower models, recent work has proposed Generative Retrieval (GR). GR models leverage a sequence of tokens using generative models, such as Transformers (Vaswani et al., 2017), and then use those tokens to retrieve Pins. GR has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (Kang and McAuley, 2018a; Li et al., 2023; Yang et al., 2024).
In this paper, we aim to develop a unified GR model that can serve multiple Pinterest recommendation surfaces (e.g., search, home feed, related Pins) by being fine-tuned on the target surface. Because GR models operate over a single token sequence, they can combine heterogeneous inputs—such as query text and user history—into one representation, enabling parameter sharing and cross-surface transfer. Moreover, the training signal is naturally cross-surface: interaction logs from different surfaces can be expressed in a common next-item (or next-token) prediction framework, allowing the model to learn from a unified mixture of surface-specific trajectories. In contrast, two-tower systems often require surface-specific query towers, objectives, and/or retrieval indices, which increases complexity and can limit generalization across surfaces.
Despite these advantages, translating a unified GR approach into a production retriever that serves multiple Pinterest surfaces introduces several practical challenges. First, it must efficiently optimize surface-specific objectives, since each surface defines success differently: search emphasizes outbound clicks that reflect offsite conversion intent, related Pins emphasizes session depth and continued exploration, and home feed emphasizes saves and discovery (clicks and downstream exploration). This is further complicated when key metrics are sparse—for example, outbound product clicks in search are relatively rare events. Second, user behavior and demographics differ by surface: upstream surfaces at the start of a session (e.g., the home feed) have a higher share of less-active users, whereas downstream surfaces (e.g., related Pins) skew toward more active users who take stronger actions. Third, user intent can evolve even within a single session, and a generative retriever must adapt its generated sequence accordingly. Fourth, due to sequential decoding, GR models typically incur higher computational cost and latency than two-tower models, which only require embedding computation and nearest-neighbor search.
Figure 2. PinRec is pretrained with the user activity sequences consisting of items (Pins) and search queries. PinRec is then fine-tuned for each surface specific data that contains user actions and impressed items (Pins).
To our knowledge, no prior work addresses all of these requirements in a single unified GR model, namely (i) surface-specific success metrics, (ii) heterogeneity in cross-surface data distributions, (iii) in-session user intent shifts, and (iv) the latency/cost constraints. Recent industrial GR systems such as TIGER and HSTU (Zhai et al., 2024; Rajput et al., 2023) are designed for a single next-token/next-item prediction objective, and in practice are typically trained against a unified engagement target (e.g., any action) rather than explicitly aligning the learning signal to surface-specific goals (e.g., outbound clicks for search, saves for home feed). As a result, fine-tuning can yield limited gains when the surface “true north” metrics differ (see Sec. 5.2.3). Separately, foundational sequence models like RecFormer and PLUM can be fine-tuned for individual applications (Hou et al., 2022; Li et al., 2023; He et al., 2025) and can better match surface-specific user/item distributions, but they are often prohibitively expensive at Pinterest scale due to their reliance on LLM backbones; moreover, they are trained for a single generic engagement target, leaving the objective-misalignment issue unchanged. Finally, existing GR approaches do not support multi-step generation at retrieval time; because autoregressive decoding is costly, they commonly generate only the next-item candidates, limiting their ability to produce a longer candidate sequence that anticipates or tracks within-session shifts in intent.
Figure 3. Generation phase of PinRec
In this paper, we present PinRec, a foundational GR model for Pinterest recommendation surfaces. We pretrain PinRec on user activity data aggregated across surfaces (Chen et al., 2025), and then fine-tune it on surface-specific data to align the model with each surface’s business goals (Fig.  2). Unlike prior GR methods, PinRec addresses four aforementioned challenges: (1) optimizing surface-specific objectives, (2) adapting to each surface’s data distribution, (3) capturing evolving user intent within a session via sequence generation, and (4) doing so with only a marginal increase in serving cost. We summarize our solutions to each challenge below.
@@
  • Outcome-Conditioned Generative Retrieval: To optimize surface-specific objectives, PinRec employs outcome-conditioned generation, enabling the optimization of multiple business metrics. For each outcome (e.g., user saves or clicks), PinRec generates a sequence of Pin tokens conditioned on that outcome and retrieves candidate Pins; the generated sequences can then be combined to balance objectives according to desired trade-offs (Fig. 3). Our offline and online evaluations show that outcome conditioning enables PinRec to adapt to the distinct business needs of each surface.
  • Fine-tuning for each surface: To adapt for each surface, we employ a pretraining–fine-tuning framework. We first pretrain the base PinRec model on user activity sequences aggregated across all Pinterest surfaces. To tailor the model to each surface’s user population and behavior patterns, we then fine-tune PinRec on surface-specific impression logs. Using this approach, we deployed PinRec in production across three surfaces. To our knowledge, this is the first paper to apply a pretraining–fine-tuning paradigm for generative retrieval across multiple recommendation surfaces at web scale.
  • Efficient Serving with Auto-regressive generation: To capture evolving user interest, PinRec generates item sequences auto-regressively. To mitigate the cost of generation, we describe optimizations that enable low-latency decoding, detail the serving infrastructure, and show how we scale the system horizontally to serve hundreds of millions of users. To our knowledge, this is the first study to disclose infrastructure optimizations for an industry-scale generative retrieval system.
@@ 2. Related Work
Sequential Recommendation Systems: Sequential recommendation systems operate on a sequences of items with which a user has interacted (Kang and McAuley, 2018a). Earlier work in academia has used recurrent neural networks (RNN) (Hidasi et al., 2016), but later work has often adopted transformers (Kang and McAuley, 2018a; Sun et al., 2019). Recent studies have harnessed the extensive knowledge of pre-trained language models (Hou et al., 2022; Li et al., 2023; He et al., 2025).
In industrial applications, a common approach is to apply sequential encoders to learn user representations (Pancha et al., 2022). The user representation can then be used in traditional recommender systems (e.g., embedding-based retrieval or ranking models) (Hansen et al., 2020; Xu et al., 2022; Xia et al., 2023; Si et al., 2024). A different line of work in sequential recommendation is generative recommendation, where the model directly generates the items to recommend. Rajput et al. (2023) uses a Transformer to generate items by producing sequences of semantic ID tokens (Lee et al., 2022; Singh et al., 2024), while Zhai et al. (2024) proposes a new architecture (HSTU). However, these early methods have largely been evaluated in single-application settings; extending them to multiple surfaces is challenging because they typically optimize only a single objective (e.g., whether an action occurs).
Very recently, Deng et al. (2025); Dai et al. (2025) proposed multi-stage training frameworks that unify retrieval and ranking within a single surface. This represents vertical unification, integrating different stages of the recommendation funnel for one surface. In contrast, we pursue horizontal unification: we fine-tune a single retrieval model to serve multiple surfaces. Both directions are important, but we expect horizontal unification to be easier to translate into product impact, since the unified model can be integrated into existing multi-stage funnels as a drop-in replacement at the retrieval stage.
Multi-Objective Generation: We discuss generative models that consider multiple objectives. While RL algorithms have shown this capability in industry (Xin et al., 2020), studies in language modeling suggest RL often needs extensive tuning for optimal generation (Ouyang et al., 2022). Recent research has introduced more streamlined and scalable techniques based on supervised learning, such as goal-conditioned methods (Ghosh et al., 2021) and prompting strategies (Xin et al., 2022). PinRec builds on these approaches to enhance scalability.
@@ 3. Problem Setup
User History. For each user $u$, we maintain their interaction history $H(u,t_{\text{max}})=\langle(i_{1},a_{1},s_{1},t_{1}),(i_{2},a_{2},s_{2},t_{2}),\ldots,(i_{m},a_{m},s_{m},t_{m})\rangle$, where each tuple $(i_{j},a_{j},s_{j},t_{j})$ represents an interaction with item $i_{j}\in\mathcal{I}$ via action $a_{j}\in\mathcal{A}$ on surface $s_{j}\in\mathcal{S}$ at timestep $t_{j}\leq t_{\text{max}}$. The sequence is chronologically ordered, $t_{1}\leq t_{2}\leq\ldots\leq t_{m}$.
Surface-specific Feed View. When user $u$ visits surface $s$ at time $t$, we observe a feed-view event $F(u,s,t)=\langle(i_{1},a_{1},t_{1}),\ldots,(i_{m},a_{m},t_{m})\rangle,$ where each tuple $(i_{j},a_{j},t_{j})$ indicates that item $i_{j}\in\mathcal{I}$ was shown to the user at $j$-th position, and the user’s action on that impression is encoded by $a_{j}\in\mathcal{A}\cup\{\text{imp}\},$ with imp denoting an impression with no explicit action taken (i.e., the item was shown but not clicked/saved/etc.). If an explicit engagement occurs (e.g., click, save), then $a_{j}\in\mathcal{A}$. The timestamp $t_{j}$ is the time of action when $a_{j}\in\mathcal{A}$; otherwise (no action), we set $t_{j}=t$.
@@ 4. PinRec
In this section, we describe the model architecture and serving infrastructure for PinRec. We start with an overview of the model, including outcome-conditioned generation. Next, we explain how items in the sequence are encoded and processed. Then we talk about how we perform pretraining on a unified data and how we fine-tune for a surface-specific data. Finally, we describe inference mode operation and provide details about our real-time serving architecture, which is powered by Nvidia Triton.
@@ 4.1. Model Architecture
Outcome-Conditioned Generation. Unlike language modeling, where the output sequence receives a single reward value (Ouyang et al., 2022), users interact with recommended items through diverse engagement types—saving, sharing, commenting, or clicking. To effectively optimize across these varied engagement metrics, generative retrieval methods must account for this behavioral diversity (Liu et al., 2024). Furthermore, practical deployment often requires specifying the target surface or context for the next recommendation, even when the user’s historical sequence spans multiple surfaces.
@@ 4.2. Input Sequence Representation
Given the heterogeneous nature of user interactions $(i_{j},a_{j},s_{j},t_{j})$ in our interaction history $H(u,t_{\text{max}})$, we need to transform each element into appropriate representations for the transformer architecture. There are two main ways to represent items in a sequence: (1) by using dense embedding vectors (Zhai et al., 2024; Kang and McAuley, 2018a), and (2) by learning a sequence of discrete IDs for each item, known as Semantic IDs (Rajput et al., 2023; Singh et al., 2024). We explored both approaches and ultimately chose to use dense embedding vectors. In our experiments, we observed that Semantic IDs frequently suffered from representational collapse—many items were assigned the same Semantic ID, as also noted in (Liu et al., 2024; Kuai et al., 2024)—and consistently produced lower evaluation metrics (see TIGER in Table 1(Yang et al., 2024). While optimizing performance with Semantic IDs remains an interesting direction for future work, our core contributions are representation-agnostic.
@@
Pin Embeddings. For Pins, we use the pre-trained OmniSage embedding (Badrinath et al., 2025), which captures visual and textual content along with the engagement via the Pin-board graph. The OmniSage embedding is then passed through 2-3 fully connected layers with ReLU activation and layer normalization, followed by $L_{2}$ normalization to produce the final Pin embedding $\mathbf{i}_{\text{pin}}$.
Query Embeddings. For search queries, we leverage the pre-trained OmniSearchSage query representation (Agarwal et al., 2024) which captures semantic query understanding. This representation is passed through a similar MLP architecture (2-3 fully connected layers with ReLU activation and layer normalization) followed by $L_{2}$ normalization to generate the final query embedding $\mathbf{i}_{\text{query}}$.
@@
Our item representations rely on embeddings and signals that are available for new items unseen during training. This enables PinRec to generalize to unseen items—such as Ads Pins and Search queries—as we show in Sec. 6.3, highlighting the strong generalization of generative retrieval models.